Use the complete the square method to solve for X.


(x)=2x² - 9x +10


1
Expert's answer
2021-06-11T16:38:14-0400
"f(x)=2x^2-9x+10"

"f(x)=0=>2x^2-9x+10=0"

"2(x^2-\dfrac{9}{2}x+5)=0"


"x^2-2(\dfrac{9}{4})x+(\dfrac{9}{4})^2-(\dfrac{9}{4})^2+5=0"


"(x-\dfrac{9}{4})^2-\dfrac{1}{16}=0"

"(x-\dfrac{9}{4}-\dfrac{1}{4})(x-\dfrac{9}{4}+\dfrac{1}{4})=0"

"(x-\dfrac{5}{2})(x-2)=0"

"x_1=2, x_2=\dfrac{5}{2}"

"\{2, \ \dfrac{5}{2}\}"